# A Wind-Driven Origin for the Firework Morphology of the Supernova Remnant Pa 30

### v1

@@ Abstract
[…]
@@ 1 Introduction
Instead, we propose that the WD wind (not the original supernova ejecta) is responsible for the Pa 30 firework morphology, as follows from RTI and KHI considerations. We expect the wind density was initially much larger than the circumstellar density (see Section 2.1). In this case, the RTI growth rate is relatively insensitive to the density contrast between the wind and its surroundings (see Section 2.2). Contrarily, the KHI growth rate scales as the square root of the ratio of the less-dense to the more-dense fluid. Thus, the KHI should be inhibited when the density contrast is large, despite the RTI-induced velocity shear. The shear itself is also minimized during this phase, because the forward shock moves out only marginally faster than the contact discontinuity, which further reduces the KHI growth rate. Moreover, the wind density declines relatively gradually (as $\propto R^{-2}$ versus $R^{-3}$ for a homologous explosion) with distance $R$ to the leading edge of the expanding wind, which further serves to maintain a high density contrast and stabilize the RTI-induced plumes against the KHI. In summary, the interaction of a dense wind with a diffuse surrounding medium should be unstable to the RTI but relatively stable to the KHI.
@@
2 Wind-driven Rayleigh-Taylor instability
@@
2.1 Model parameters
Observations of Pa 30 place the WD wind speed at ${\sim}10^{4}$ km s-1 and its mass loss rate at ${\simeq}10^{-6}{-}10^{-7}M_{\odot}$ yr-1 (Lykou et al., 2023). These WD wind properties lead to comparable initial densities across the CD that would not be conducive to RTI-induced filament formation (see below and Section 2.2). However, the natal WD merger remnant likely had a lower wind speed $V_{\rm w}\lesssim 3\times 10^{3}$ km s-1, and a higher mass-loss rate $\dot{M}\simeq 10^{-5}M_{\odot}$ yr-1 (Shen & Schwab, 2017); we adopt these as fiducial initial values that applied near the supernova epoch. Notably, the initial mass-loss rate may have been even larger owing to the likely ONe-CO progenitor (Oskinova et al., 2020), the very low-energy nature of the explosion (Shao et al., 2025) plus substantial fallback of the initial ejecta, and the corresponding possibility of an initially hyper-accreting disc launching the wind (instead of the delayed radioactive decays as described in Shen & Schwab 2017); we revisit this possibility at the end of Section 4. Scaling the radius $R$ to the WD surface at $0.01R_{\odot}$, the wind density profile is
@@ 2.1 Model parameters
We assume the ambient medium is homogeneous and confined to $\sim$}sub-pc scales (i.e., interior to the ISM), such that its density does not decline much with distance from the central WD driving the wind. [^4] Section 3 discusses the importance of the spatial profile of the ambient medium. In the limit of infinite density contrast, any small-scale inhomogeneities/clumps in the ambient medium will excite perturbations that evolve largely independently of perturbations to the CD/shocked wind. In this limit, an analysis of the perturbations imparted to the shocked ambient gas is more tractable (than is the case for the shocked wind) because they are separable in $\log(t)$ and the spatial self-similar variable (e.g., Ryu & Vishniac 1987; Goodman 1990); it is straightforward to show that they are stable with the eigenmodes decaying with time approximately as $\propto t^{-3}$. Therefore, inhomogeneities in the shocked ambient medium will largely produce damped acoustic oscillations (cf. Coughlin & Zrake 2022) that do not significantly modify the shocked wind, although they could still seed the RTI if they are large.
2.2 Stability and growth
@@
Note that unlike the standard RTI, the dispersion relation above depends on the density contrast. This arises because the pressure gradient that drives the RTI is established by the CD deceleration, which itself depends on the density ratio. [^5] We discuss this feature and its implications in Section 4. Although the density contrast acts to slow the RTI growth, $\tau_{\rm RT}\propto\left(\rho_{\rm w}/\rho_{\rm a}\right)^{1/4}$, it more effectively slows the KHI growth, $\tau_{\rm KH}\propto\left(\rho_{\rm w}/\rho_{\rm a}\right)^{1/2}$.
2.3 Filament properties
The RTI should stop operating roughly when the wind and ambient densities are equal; [^6] Strictly speaking it is also when there is no longer a strong pressure gradient in the shocked wind relative to the shocked ambient medium, which should occur roughly simultaneously with when the densities become comparable, as at this point the FS decelerates more substantially and the solution enters the self-similar phase described in Weaver et al. (1977). that is, when $R_{\rm c}/R_{\rm i}=10{-}100$, or $R_{\rm c}=10^{16-17}$ cm (see Section 2.1). Reaching this equality takes just $\sim$}1–10 yr for a wind speed of $10^{3{-}4}$ km s-1. The RTI growth timescale of the filaments is shortest for the smallest-scale perturbations (i.e., largest $k$; see Section 2.2). Consequently, the filaments should be highly elongated with a length that is limited by the timescale over which the RTI operates.
@@ 2.3 Filament properties
where the numerical factor of $0.46$ arises from the assumption that the FS shell is gas-pressure dominated (it is $0.36$ for a radiation-dominated gas). This is faster than the initial speed of the FS, which is ${\sim}0.1V_{\rm c}$ (again relative to the CD; see Equation 4), implying that the RTI plumes could overtake the FS (as seen in simulations; Duffell et al. 2024; Blondin & Ellison 2001). That said, Equation (8) is likely a substantial overestimate because it crudely assumes uniform pressure within the plumes, despite their implied velocity being supersonic. [^7] Due to the large density contrast, the shocked wind sound speed is $\sqrt{\rho_{\rm a}/\rho_{\rm w}}V_{c}\ll 0.46V_{c}$; thus, the plumes are supersonic. Therefore, the filaments likely propagate downstream only marginally faster than the CD, whose lab-frame speed $V_{\rm c}$ is of order the natal WD wind speed (${\sim}{\rm few}\times 10^{3}$ km s-1; see Section 2.1) and is consistent with the observed speed of the filaments (Cunningham et al., 2024).
@@
From the considerations above, the filaments should expand primarily into the shocked ambient medium; grow predominantly in the radial direction, because their large densities and low shear inhibit efficient KHI growth; and decline in density as ${\sim}1/r$, because they are pressure confined. Therefore, we expect the RTI-induced filaments to have large aspect ratios and to extend approximately radially outward from the system center. If the wind density was ${\sim}10^{-19}$ g cm-3 at $R_{\rm i}=10^{15}$ cm (see Section 2.1), then after radially expanding by a factor of ${\sim}100$ (at which point the RTI should cease), their densities would fall to ${\sim}10^{-21}$ g cm-3. This corresponds to a number density of ${\sim}{\rm few}\times 100$ cm-3, which is consistent with those observed of ${\sim}200{-}1000$ cm-3 based on [S II] emission lines (Fesen et al., 2023).
@@
The initial temperature of the shocked (assumed gas-pressure-dominated) wind is $T\simeq 10^{5{-}6}$ K, which arises from a density contrast of $\rho_{\rm w}/\rho_{\rm a}\simeq 10^{3}$ and a wind speed of ${\sim}{\rm few}\times 10^{3}$ km s-1 (see Section 2.1). If the filaments grow adiabatically, then their temperature scales as $\rho^{2/3}$ and falls to ${\sim}10^{3{-}4}$ K by the time they expand by a factor of ${\sim}10^{2{-}3}$ to reach pc-scale lengths (in line with inferences in Cunningham et al. 2024). Therefore, we expect the filaments to be optically bright but largely devoid of hydrogen (because they originate from a WD-driven wind at early epochs), which is consistent with observations.
3 Discussion
@@
Pa 30 is seemingly the only known SNR with a firework-like morphology, which begs the question of the origin of its uniqueness. As noted in the introduction, its high-speed WD wind offers a suggestive clue. Our model incorporates an early-phase WD wind, driven by either radioactive decay of freshly synthesized elements (Shen & Schwab, 2017) or supercritical disc accretion (see Section 4). We propose that the RTI generates tendrils of high-density, shocked wind material that propagate into the low-density, shocked ambient medium. The high mass-loss rate from the natal WD wind implies that its density remains large relative to the CSM out to large distances. Crucial to our proposed RTI filament formation mechanism, the system thus maintains a high density contrast without quickly transitioning to the self-similar solution described in Weaver et al. (1977); as the nearly isobaric shocked wind would be RTI-stable in this wind-driven analog of the Sedov-Taylor blast wave. Presumably, other supernovae also go through a phase of high density contrast (e.g., Björnsson 2025), yet they do not display a firework morphology. In some cases, this phase is likely too short-lived for filaments to develop. For instance, type Ibc and type II supernovae enshrouded in a substantial gaseous reservoir (e.g., from late-stage stellar winds and eruptions) likely lack the necessary large density contrast extending out to large distances.
@@
The Pa 30 filaments display faint [O III] emission (Cunningham et al., 2024) but the WD and its wind are currently oxygen-rich (Oskinova et al., 2020). If the filaments are made of WD wind material, then why are these oxygen levels so different? As noted above, during the filament formation epoch, the wind-driving mechanism was probably related to the radioactive decay of newly synthesized elements following the merger-induced deflagration, which failed to produce a detonation (Shen & Schwab, 2017), or to a hypercritical phase of accretion (Section 4). Therefore, during the short time that the filaments formed and elongated ($\sim$1–10 years; Section 2.3), the natal wind was likely polluted by heavy elements; thus, explaining their low oxygen content.
@@
Our simple hydrodynamic simulations (see Figure 1) used a single-mode perturbation to seed the RTI, leading to spikes of equal length and spacing. A more stochastic distribution of fluctuations within the ejecta would produce a more varied distribution of tendril lengths and widths. Ambient medium fluctuations, if sufficiently strong (see end of Section 2.1), could also contribute to tendril variety. Given that the RTI growth rate likely scales with spherical harmonic as $\sigma_{\rm RT}\propto\ell^{1/2}$ at sufficiently high $\ell$ (i.e., as the standard RTI analysis predicts; see Equation 5), one might expect narrower filaments to form first but then truncate (due to the onset of wider filaments sapping the mass supply), leading to a small-scale clumpy structure. However, this outcome depends on complex couplings involving the non-linear instability growth, the evolving RS shell width, and extensions to the standard RTI analysis. Imminent imaging of Pa 30 by the James Webb Space Telescope and the Hubble Space Telescope will provide further insight into the coherence of the filaments.
4 Summary and Conclusions
@@
We proposed an explanation for the firework-like morphology of the Pa 30 SNR. The natal WD drives a high-density wind into the CSM, whose density is low and declines gradually compared to a homologous explosion (Section 2.1). The large density contrast at the CD enables the RTI to grow with an approximate $e$-folding timescale of $\tau_{\rm RT}\sim\left(R_{\rm i}/V_{\rm w}\right)\times\left(\rho_{\rm a}/\rho_{\rm w}\right)^{1/4}\times k^{-1/2}$, with $k$ in units of $R_{\rm i}^{-1}$ (Section 2.2). A continuous supply of shocked wind material gets channeled into the RTI spikes as they elongate toward the FS, all the while draining the reservoir of shocked wind material in the RS shell. [^9] The compounding effect of low-density bubbles (made of shocked ambient gas) moving into the RS shell further lowers its density. Consequently, the CD reaches density parity after just $\sim$}1–10 yr and the RTI ceases, which truncates the spikes (Section 2.3). The spikes remain intact because their large density contrast with the shocked CSM and low shear velocity—a consequence of the FS moving at nearly the same speed as the CD—prevent their destruction by KHI (Section 2.3).
@@
At a high level, our model suggests geometric, kinematic, and thermodynamic properties of the filaments that are consistent with observations (Sections 2.3 and 3). An important aspect of our model is that the dense wind of the natal WD remnant—not the preceding type Iax supernova ejecta—primarily dictates the firework morphology of Pa 30, in line with radio upper limits that imply a weak explosion (Shao et al., 2025). Similar morphologies could be present in other type Iax systems, especially if a prominent feature of their evolution is a dense wind encountering a rarefied CSM (as suggested by, e.g., Camacho-Neves et al. 2023; Callan et al. 2024; Kwok et al. 2025; Schwab et al. 2025).
@@
We also expect the dispersion relation to change qualitatively once the perturbation length scale is larger than the separation between the RS and CD. [^10] If one accounts for the finite vertical extent of the upper fluid, $H$, the standard RTI dispersion relation becomes $\sigma_{\rm RTI}^{2}=\frac{\rho_{\rm d}-\rho_{\rm u}}{\rho_{\rm u}\tanh(|k|H)+\rho_{\rm d}}|k|g$, implying that the growth rate is suppressed once $|k|H\lesssim\rho_{\rm d}/\rho_{\rm u}$; since $H\propto\sqrt{\rho_{\rm d}/\rho_{\rm a}}$, this translates to $k\lesssim 1$ when $k$ is measured in units of $R_{\rm i}^{-1}$. If the initial encounter between the wind and the ambient medium takes place at $R_{\rm i}=10^{15}$ cm with a density contrast of $10^{3}$, then the initial width of the shocked wind layer would be ${\sim}0.0484\times 10^{-3/2}R_{\rm i}\simeq 10^{12}$ cm. If the maximum transverse size of RTI-unstable modes were limited to this length scale (to within a factor of ${\sim}\pi$), the filament aspect ratio would decline from $1/1000$ with time, implying they would be unobservable (not to mention much narrower than their observed aspect ratio of ${\sim}0.01{-}0.1$; Cunningham et al. 2024). However, it is unclear if $R_{\rm i}$ alone sets the width of the filaments, or if their widths expand as the CD radius and RS shell width expand. Unfortunately, the methodology in Section 2.2 cannot differentiate between these possibilities—the WKB approximation implicitly assumes that the growth timescale is much shorter than $R_{\rm i}/V_{\rm w}$. We intend to explore the nature of the RTI during this phase both numerically and analytically in future work.
@@
Finally and as previously noted, radio observations of Pa 30 imply that the 1181 AD type Iax supernova was of exceptionally low energy (Shao et al., 2025), and the spectroscopically inferred elemental abundances of the central star suggest that the progenitor consisted of the merger of a high-mass ONe WD (mass $\gtrsim 1~M_{\odot}$) and a lower-mass CO WD (Oskinova et al., 2020). The merger therefore likely proceeded through the tidal disruption of the CO WD (e.g., Kashyap et al. 2018), resulting in the formation of an accretion flow around the ONe WD disruptor and leaving much of the explosion ejecta bound to the system to fuel additional, later-time fallback. Therefore, instead of the initial wind being launched by the delayed radioactive decay of heavy elements (Shen & Schwab, 2017), it may have been driven by a period of sustained supercritical accretion with mass loss rates presumably much larger than assumed in Section 2.1. Similar physical conditions (an overdense wind impacting a medium) may be present in, e.g., the collapsar model of long gamma-ray bursts (MacFadyen & Woosley, 1999), delayed circularization and black-hole disc formation in failed supernovae (Quataert et al., 2019), hyperaccreting neutron stars/ultra-luminous X-ray sources (e.g., Strohmayer & Mushotzky 2003), and the winds from tidal disruption events (e.g., Miller et al. 2015), to which it would be interesting to apply our model—both in the context of the wind-CSM dynamics and the possibility of filament formation.

### v2

@@ Abstract
analytical mathematics (38) — hydrodynamics (1963) — shocks (2086) — type Ia supernovae (1728)
@@ I Introduction
Instead, we propose that the WD wind (not the original supernova ejecta) is responsible for the Pa 30 firework morphology, as follows from RTI and KHI considerations. We expect the wind density was initially much larger than the circumstellar density (see Section II.1). In this case, the RTI growth rate is relatively insensitive to the density contrast between the wind and its surroundings (see Section II.2). Contrarily, the KHI growth rate scales as the square root of the ratio of the less-dense to the more-dense fluid. Thus, the KHI should be inhibited when the density contrast is large, despite the RTI-induced velocity shear. The shear itself is also minimized during this phase, because the forward shock moves out only marginally faster than the contact discontinuity, which further reduces the KHI growth rate. Moreover, the wind density declines relatively gradually (as $\propto R^{-2}$ versus $R^{-3}$ for a homologous explosion) with distance $R$ to the leading edge of the expanding wind, which further serves to maintain a high density contrast and stabilize the RTI-induced plumes against the KHI. In summary, the interaction of a dense wind with a diffuse surrounding medium should be unstable to the RTI but relatively stable to the KHI.
@@
II Wind-driven Rayleigh-Taylor instability
@@
II.1 Model parameters
Observations of Pa 30 place the WD wind speed at ${\sim}10^{4}$ km s-1 and its mass loss rate at ${\simeq}10^{-6}{-}10^{-7}M_{\odot}$ yr-1 (Lykou et al., 2023). These WD wind properties lead to comparable initial densities across the CD that would not be conducive to RTI-induced filament formation (see below and Section II.2). However, the natal WD merger remnant likely had a lower wind speed $V_{\rm w}\lesssim 3\times 10^{3}$ km s-1, and a higher mass-loss rate $\dot{M}\simeq 10^{-5}M_{\odot}$ yr-1 (Shen & Schwab, 2017); we adopt these as fiducial initial values that applied near the supernova epoch. Notably, the initial mass-loss rate may have been even larger owing to the likely ONe-CO progenitor (Oskinova et al., 2020), the very low-energy nature of the explosion (Shao et al., 2025) plus substantial fallback of the initial ejecta, and the corresponding possibility of an initially hyper-accreting disc launching the wind (instead of the delayed radioactive decays as described in Shen & Schwab 2017); we revisit this possibility at the end of Section IV. Scaling the radius $R$ to the WD surface at $0.01R_{\odot}$, the wind density profile is
@@ II.1 Model parameters
We assume the ambient medium is homogeneous and confined to $\sim$}sub-pc scales (i.e., interior to the ISM), such that its density does not decline much with distance from the central WD driving the wind. [^4] Section III discusses the importance of the spatial profile of the ambient medium. In the limit of infinite density contrast, any small-scale inhomogeneities/clumps in the ambient medium will excite perturbations that evolve largely independently of perturbations to the CD/shocked wind. In this limit, an analysis of the perturbations imparted to the shocked ambient gas is more tractable (than is the case for the shocked wind) because they are separable in $\log(t)$ and the spatial self-similar variable (e.g., Ryu & Vishniac 1987; Goodman 1990); it is straightforward to show that they are stable with the eigenmodes decaying with time approximately as $\propto t^{-3}$. Therefore, inhomogeneities in the shocked ambient medium will largely produce damped acoustic oscillations (cf. Coughlin & Zrake 2022) that do not significantly modify the shocked wind, although they could still seed the RTI if they are large.
II.2 Stability and growth
@@
Note that unlike the standard RTI, the dispersion relation above depends on the density contrast. This arises because the pressure gradient that drives the RTI is established by the CD deceleration, which itself depends on the density ratio. [^5] We discuss this feature and its implications in Section IV. Although the density contrast acts to slow the RTI growth, $\tau_{\rm RT}\propto\left(\rho_{\rm w}/\rho_{\rm a}\right)^{1/4}$, it more effectively slows the KHI growth, $\tau_{\rm KH}\propto\left(\rho_{\rm w}/\rho_{\rm a}\right)^{1/2}$.
II.3 Filament properties
The RTI should stop operating roughly when the wind and ambient densities are equal; [^6] Strictly speaking it is also when there is no longer a strong pressure gradient in the shocked wind relative to the shocked ambient medium, which should occur roughly simultaneously with when the densities become comparable, as at this point the FS decelerates more substantially and the solution enters the self-similar phase described in Weaver et al. (1977). that is, when $R_{\rm c}/R_{\rm i}=10{-}100$, or $R_{\rm c}=10^{16-17}$ cm (see Section II.1). Reaching this equality takes just $\sim$}1–10 yr for a wind speed of $10^{3{-}4}$ km s-1. The RTI growth timescale of the filaments is shortest for the smallest-scale perturbations (i.e., largest $k$; see Section II.2). Consequently, the filaments should be highly elongated with a length that is limited by the timescale over which the RTI operates.
@@ II.3 Filament properties
where the numerical factor of $0.46$ arises from the assumption that the FS shell is gas-pressure dominated (it is $0.36$ for a radiation-dominated gas). This is faster than the initial speed of the FS, which is ${\sim}0.1V_{\rm c}$ (again relative to the CD; see Equation 4), implying that the RTI plumes could overtake the FS (as seen in simulations; Duffell et al. 2024; Blondin & Ellison 2001). That said, Equation (8) is likely a substantial overestimate because it crudely assumes uniform pressure within the plumes, despite their implied velocity being supersonic. [^7] Due to the large density contrast, the shocked wind sound speed is $\sqrt{\rho_{\rm a}/\rho_{\rm w}}V_{c}\ll 0.46V_{c}$; thus, the plumes are supersonic. Therefore, the filaments likely propagate downstream only marginally faster than the CD, whose lab-frame speed $V_{\rm c}$ is of order the natal WD wind speed (${\sim}{\rm few}\times 10^{3}$ km s-1; see Section II.1) and is consistent with the observed speed of the filaments (Cunningham et al., 2024).
@@
From the considerations above, the filaments should expand primarily into the shocked ambient medium; grow predominantly in the radial direction, because their large densities and low shear inhibit efficient KHI growth; and decline in density as ${\sim}1/r$, because they are pressure confined. Therefore, we expect the RTI-induced filaments to have large aspect ratios and to extend approximately radially outward from the system center. If the wind density was ${\sim}10^{-19}$ g cm-3 at $R_{\rm i}=10^{15}$ cm (see Section II.1), then after radially expanding by a factor of ${\sim}100$ (at which point the RTI should cease), their densities would fall to ${\sim}10^{-21}$ g cm-3. This corresponds to a number density of ${\sim}{\rm few}\times 100$ cm-3, which is consistent with those observed of ${\sim}200{-}1000$ cm-3 based on [S II] emission lines (Fesen et al., 2023).
@@
The initial temperature of the shocked (assumed gas-pressure-dominated) wind is $T\simeq 10^{5{-}6}$ K, which arises from a density contrast of $\rho_{\rm w}/\rho_{\rm a}\simeq 10^{3}$ and a wind speed of ${\sim}{\rm few}\times 10^{3}$ km s-1 (see Section II.1). If the filaments grow adiabatically, then their temperature scales as $\rho^{2/3}$ and falls to ${\sim}10^{3{-}4}$ K by the time they expand by a factor of ${\sim}10^{2{-}3}$ to reach pc-scale lengths (in line with inferences in Cunningham et al. 2024). Therefore, we expect the filaments to be optically bright but largely devoid of hydrogen (because they originate from a WD-driven wind at early epochs), which is consistent with observations.
III Discussion
@@
Pa 30 is seemingly the only known SNR with a firework-like morphology, which begs the question of the origin of its uniqueness. As noted in the introduction, its high-speed WD wind offers a suggestive clue. Our model incorporates an early-phase WD wind, driven by either radioactive decay of freshly synthesized elements (Shen & Schwab, 2017) or supercritical disc accretion (see Section IV). We propose that the RTI generates tendrils of high-density, shocked wind material that propagate into the low-density, shocked ambient medium. The high mass-loss rate from the natal WD wind implies that its density remains large relative to the CSM out to large distances. Crucial to our proposed RTI filament formation mechanism, the system thus maintains a high density contrast without quickly transitioning to the self-similar solution described in Weaver et al. (1977); as the nearly isobaric shocked wind would be RTI-stable in this wind-driven analog of the Sedov-Taylor blast wave. Presumably, other supernovae also go through a phase of high density contrast (e.g., Björnsson 2025), yet they do not display a firework morphology. In some cases, this phase is likely too short-lived for filaments to develop. For instance, type Ibc and type II supernovae enshrouded in a substantial gaseous reservoir (e.g., from late-stage stellar winds and eruptions) likely lack the necessary large density contrast extending out to large distances.
@@
The Pa 30 filaments display faint [O III] emission (Cunningham et al., 2024) but the WD and its wind are currently oxygen-rich (Oskinova et al., 2020). If the filaments are made of WD wind material, then why are these oxygen levels so different? As noted above, during the filament formation epoch, the wind-driving mechanism was probably related to the radioactive decay of newly synthesized elements following the merger-induced deflagration, which failed to produce a detonation (Shen & Schwab, 2017), or to a hypercritical phase of accretion (Section IV). Therefore, during the short time that the filaments formed and elongated ($\sim$1–10 years; Section II.3), the natal wind was likely polluted by heavy elements; thus, explaining their low oxygen content.
@@
Our simple hydrodynamic simulations (see Figure 1) used a single-mode perturbation to seed the RTI, leading to spikes of equal length and spacing. A more stochastic distribution of fluctuations within the ejecta would produce a more varied distribution of tendril lengths and widths. Ambient medium fluctuations, if sufficiently strong (see end of Section II.1), could also contribute to tendril variety. Given that the RTI growth rate likely scales with spherical harmonic as $\sigma_{\rm RT}\propto\ell^{1/2}$ at sufficiently high $\ell$ (i.e., as the standard RTI analysis predicts; see Equation 5), one might expect narrower filaments to form first but then truncate (due to the onset of wider filaments sapping the mass supply), leading to a small-scale clumpy structure. However, this outcome depends on complex couplings involving the non-linear instability growth, the evolving RS shell width, and extensions to the standard RTI analysis. Imminent imaging of Pa 30 by the James Webb Space Telescope and the Hubble Space Telescope will provide further insight into the coherence of the filaments.
IV Summary and Conclusions
@@
We proposed an explanation for the firework-like morphology of the Pa 30 SNR. The natal WD drives a high-density wind into the CSM, whose density is low and declines gradually compared to a homologous explosion (Section II.1). The large density contrast at the CD enables the RTI to grow with an approximate $e$-folding timescale of $\tau_{\rm RT}\sim\left(R_{\rm i}/V_{\rm w}\right)\times\left(\rho_{\rm a}/\rho_{\rm w}\right)^{1/4}\times k^{-1/2}$, with $k$ in units of $R_{\rm i}^{-1}$ (Section II.2). A continuous supply of shocked wind material gets channeled into the RTI spikes as they elongate toward the FS, all the while draining the reservoir of shocked wind material in the RS shell. [^9] The compounding effect of low-density bubbles (made of shocked ambient gas) moving into the RS shell further lowers its density. Consequently, the CD reaches density parity after just $\sim$}1–10 yr and the RTI ceases, which truncates the spikes (Section II.3). The spikes remain intact because their large density contrast with the shocked CSM and low shear velocity—a consequence of the FS moving at nearly the same speed as the CD—prevent their destruction by KHI (Section II.3).
@@
At a high level, our model suggests geometric, kinematic, and thermodynamic properties of the filaments that are consistent with observations (Sections II.3 and III). An important aspect of our model is that the dense wind of the natal WD remnant—not the preceding type Iax supernova ejecta—primarily dictates the firework morphology of Pa 30, in line with radio upper limits that imply a weak explosion (Shao et al., 2025). Similar morphologies could be present in other type Iax systems, especially if a prominent feature of their evolution is a dense wind encountering a rarefied CSM (as suggested by, e.g., Camacho-Neves et al. 2023; Callan et al. 2024; Kwok et al. 2025; Schwab et al. 2025).
@@
Figure 2: Photographs of the “Kingfish” high-altitude nuclear test (LA-UR-25-31831). The left image was taken ${\sim}40$ ms post-detonation, illustrating the formation of tendril-like filaments in the interior of the explosion. The right image was taken ${\sim}256$ ms post-detonation, showing the transformation of the firework-like filaments into a more cauliflower-like morphology reminiscent of standard supernova remnants. This terrestrial experiment suggests that homologous explosions may also pass through a Pa 30-like evolutionary phase that is (compared to wind-driven explosions) short-lived.
@@
We also expect the dispersion relation to change qualitatively once the perturbation length scale is larger than the separation between the RS and CD. [^10] If one accounts for the finite vertical extent of the upper fluid, $H$, the standard RTI dispersion relation becomes $\sigma_{\rm RTI}^{2}=\frac{\rho_{\rm d}-\rho_{\rm u}}{\rho_{\rm u}\tanh(|k|H)+\rho_{\rm d}}|k|g$, implying that the growth rate is suppressed once $|k|H\lesssim\rho_{\rm d}/\rho_{\rm u}$; since $H\propto\sqrt{\rho_{\rm d}/\rho_{\rm a}}$, this translates to $k\lesssim 1$ when $k$ is measured in units of $R_{\rm i}^{-1}$. If the initial encounter between the wind and the ambient medium takes place at $R_{\rm i}=10^{15}$ cm with a density contrast of $10^{3}$, then the initial width of the shocked wind layer would be ${\sim}0.0484\times 10^{-3/2}R_{\rm i}\simeq 10^{12}$ cm. If the maximum transverse size of RTI-unstable modes were limited to this length scale (to within a factor of ${\sim}\pi$), the filament aspect ratio would decline from $1/1000$ with time, implying they would be unobservable (not to mention much narrower than their observed aspect ratio of ${\sim}0.01{-}0.1$; Cunningham et al. 2024). However, it is unclear if $R_{\rm i}$ alone sets the width of the filaments, or if their widths expand as the CD radius and RS shell width expand. Unfortunately, the methodology in Section II.2 cannot differentiate between these possibilities—the WKB approximation implicitly assumes that the growth timescale is much shorter than $R_{\rm i}/V_{\rm w}$. We intend to explore the nature of the RTI during this phase both numerically and analytically in future work.
@@
As previously noted, radio observations of Pa 30 imply that the 1181 AD type Iax supernova was of exceptionally low energy (Shao et al., 2025), and the spectroscopically inferred elemental abundances of the central star suggest that the progenitor consisted of the merger of a high-mass ONe WD (mass $\gtrsim 1~M_{\odot}$) and a lower-mass CO WD (Oskinova et al., 2020). The merger therefore likely proceeded through the tidal disruption of the CO WD (e.g., Kashyap et al. 2018), resulting in the formation of an accretion flow around the ONe WD disruptor and leaving much of the explosion ejecta bound to the system to fuel additional, later-time fallback. Therefore, instead of the initial wind being launched by the delayed radioactive decay of heavy elements (Shen & Schwab, 2017), it may have been driven by a period of sustained supercritical accretion with mass loss rates presumably much larger than assumed in Section II.1. Similar physical conditions (an overdense wind impacting a medium) may be present in, e.g., the collapsar model of long gamma-ray bursts (MacFadyen & Woosley, 1999), delayed circularization and black-hole disc formation in failed supernovae (Quataert et al., 2019), hyperaccreting neutron stars/ultra-luminous X-ray sources (e.g., Strohmayer & Mushotzky 2003), and the winds from tidal disruption events (e.g., Miller et al. 2015), to which it would be interesting to apply our model—both in the context of the wind-CSM dynamics and the possibility of filament formation.
Finally, while Pa 30-like supernova remnants are seemingly rare, terrestrial explosions provide another route for testing the hypothesis that tendril-like structures can form during the early phases of explosions (as suggested by the generalized dispersion relation in Section III). Figure 2 gives one such example from the “Kingfish” high-altitude nuclear test conducted by the United States in 1962: the left panel, which is a photograph of the explosion taken ${\sim}40$ ms following the initial detonation, illustrates clear evidence of radial filaments that developed. The right panel shows a photograph of the same explosion, but ${\sim}256$ ms post-detonation, showing the filaments starting to evolve into more of a cauliflower-like structure. This result suggests that homologous explosions (i.e., more typical supernovae) may pass through a morphological phase more similar to that of Pa 30, but with a temporal duration that is more short-lived than that of a wind-driven explosion.